# Photography Perspective Composition: Towards Aesthetic Perspective Recommendation

### v2

@@ 4.2 Investigation for Perspective Quality Assessment (PQA)
Implementation and Evaluation Metric. Details can be found in Appendix C.
Main Results. [Quantitative Result] To investigate the impact of training data volume in stage 1, we conducted experiments with varying numbers of video pairs, as detailed in Tab 2a.
@@
Our results reveal that while performance generally improves with increasing sample size, it plateaus at approximately 100 samples, suggesting that this quantity provides sufficient diversity for model performance. [Qualitative Results] Fig. 10 demonstrates the basic effects of PQA. As shown in Fig. 4a, while there is a notable improvement in composition, the video quality remains low. Fig. 4b exhibits acceptable levels of both VQ and MQ, but shows minimal compositional differences, resulting in a lower CA score. Notably, PQA assigns lower ratings to cases where the perspective remains static or too intense, as illustrated in the last two examples.
@@ 5 Conclusion and Limitation
Conclusion. In this work, we addressed the limitations of previous photography composition methods by introducing photography perspective composition (PPC), a novel paradigm that extends beyond 2D cropping to achieve 3D recomposition. Our approach is inspired by real-world street photography practices where photographers use perspective adjustment to establish better relative relationships between subjects. To overcome the challenges of implementing PPC, particularly the lack of suitable datasets and unclear assessment criteria, we made three significant contributions.
Figure 11: Diverse data presents instability.
We developed a framework for automatically constructing a PPC dataset from expert photographs, created a system for generating perspective transformation videos that guide users from suboptimal to optimal views, and introduced the perspective quality assessment (PQA) model that evaluates both video quality and compositional aesthetics. We hope this work opens up new possibilities in computational photography and inspires further research in perspective-aware composition.
@@
(3) Unstable Training. Despite the strong scene diversity provided by numerous expert photography images, we observed that model outputs become unstable as the training data volume increases. As shown in Tab. 2c and Fig. 4a, while both accuracy and quality initially improve with increasing data volume, performance deteriorates when the training set size grows further. We hypothesize that this challenge lies in maintaining the desired model behavior to ensure proper perspective rather than deviating into unintended random behaviors. This phenomenon bears similarities to what was described in IC-Light zhang2025scaling , and we plan to incorporate their training methodology to investigate whether it can enhance training stability in our future work.
@@ Appendix A Prompt Instructions
Appendix B Model Setting of PQA
Following VideoAlign liu2025improving , we adopt a Bradley-Terry model with ties (BTT)rao1967ties as our baseline model, which extends the traditional Bradley-Terry frameworkbradley1952rank to handle tied preferences. For a given prompt $s$ and two generated videos $v_{1}$ and $v_{2}$, the model defines the probabilities as:
$P(v_{1}=v_{2}|s)$ $=\frac{(\theta^{2}-1)e^{r(v_{1},s)}e^{r(v_{2},s)}}{\left(e^{r(v_{1},s)}+\theta e^{r(v_{2},s)}\right)\left(\theta e^{r(v_{1},s)}+e^{r(v_{2},s)}\right)},$ (3)
$P(v_{1}\succ v_{2}|s)$ $=\frac{e^{r(v_{1},s)}}{e^{r(v_{2},s)}+\theta e^{r(v_{1},s)}},$ (4)
$P(v_{2}\succ v_{1}|s)$ $=\frac{e^{r(v_{2},s)}}{\theta e^{r(v_{1},s)}+e^{r(v_{2},s)}},$ (5)
where the tie parameter $\theta$ is empirically set to $5.0$. We empirically set $\theta=5.0$ and optimize the BTT model by minimizing the negative log-likelihood (NLL):
$\mathcal{L}_{\text{BTT}}=-\mathbb{E}_{(s,v_{1},v_{2})\sim\mathcal{D}}\Big[$ $\mathds{1}(v_{1}\succ v_{2})\log P(v_{1}\succ v_{2}|s)+$ (6)
$\mathds{1}(v_{2}\succ v_{1})\log P(v_{2}\succ v_{1}|s)+$
$\mathds{1}(v_{1}=v_{2})\log P(v_{1}=v_{2}|s)\Big].$
Appendix C Implementation Details of PQA
We utilize Qwen2-VL-2B wang2024qwen2 as the backbone for PQA and train it with BTT loss. To fine-tune the model, LoRA hu2021lora is applied to update all linear layers in the language model, while the vision encoder’s parameters are fully optimized. The training process is conducted with a batch of 32 and a learning rate of $2\times 10^{-6}$, with the model trained over two epochs. This setup requires approximately 50 NVIDIA H20 GPU hours. Several observations were made during training. We sample videos at 1 fps, with a resolution of $448\times 448$ pixels during the training process.
Evaluation Metric. Following previous works, we adopt accuracy as the metric for each dimension.
Appendix D Grade Assessment in Data Filering
We implement a comprehensive grading assessment that converts model-generated scores into standardized grades scores. It employs a five-tier grading scale (A to E) with corresponding numerical ranges, as shown in Tab. 5.
Table 5: Grade assessment in data filtering.
A Excellent $\geq 5.0$ 95 (90-100) Superior quality
B Good 0.0 to 4.9 85 (80-89) Above-average quality
C Satisfactory -5.0 to -0.1 75 (70-79) Acceptable quality
D Marginal -15.0 to -5.1 65 (60-69) Below-average quality
E Unsatisfactory <15.0<-15.0 50 (<60<60) Inadequate quality
Appendix E Metric Design in PPC
For video quality assessment, we primarily adopt the evaluation metrics from Vbench2.0 huang2024vbench I2V benchmarks, which include I2V subject, I2V background subject consistency, background consistency, motion smoothness, dynamic degree, aesthetic quality, and image quality. For perspective accuracy evaluation, we employed both video distance and image similarity metrics. The video distance was measured using camera motion matching (CMM) and Fréchet Video Distance (FVD) unterthiner2019fvd , while image similarity was assessed using CLIP-F radford2021learning , PSNR, SSIM, and LPIPS. Notably, we modified the camera motion section of the original vbench huang2024vbench , changing it to camera motion matching. As we found that the camera motion detection model in vbench did not work well, especially for small angle changes. Instead, we now pass both the prediction and ground truth through the camera motion detection and then calculate the accuracy by matching these two detection results." As for the human performance score, its purpose is to simulate human preferences in scoring. This is conducted through PQA, which we established in Sec. 3.4, and includes three components: visual quality (VQ), motion quality (MQ), and composition aesthetic (CA).

### v3

@@ 4.2 Investigation for Perspective Quality Assessment (PQA)
Implementation and Evaluation Metric. We utilize Qwen2-VL-2B wang2024qwen2 as the backbone for PQA and train it with BTT loss. To fine-tune the model, LoRA hu2021lora is applied to update all linear layers in the language model, while the vision encoder’s parameters are fully optimized. The training process is conducted with a batch of 32 and a learning rate of $2\times 10^{-6}$, with the model trained over two epochs. This setup requires approximately 50 NVIDIA H20 GPU hours. Several observations were made during training. We sample videos at 1 fps, with a resolution of $448\times 448$ pixels during the training process. Following previous works liu2025improving , we adopt accuracy as the metric for each dimension.
@@
Main Results. [Quantitative Result] To investigate the impact of training data volume in stage 1, we conducted experiments with varying numbers of video pairs, as detailed in Tab 3a. Our results reveal that while performance generally improves with increasing sample size, it plateaus at approximately 100 samples, suggesting that this quantity provides sufficient diversity for model performance. [Qualitative Results] Fig. 10 demonstrates the basic effects of PQA. As shown in Fig. 4a, while there is a notable improvement in composition, the video quality remains low. Fig. 4b exhibits acceptable levels of both VQ and MQ, but shows minimal compositional differences, resulting in a lower CA score. Notably, PQA assigns lower ratings to cases where the perspective remains static or too intense, as illustrated in the last two examples.
@@ 5 Conclusion and Limitation
Conclusion. In this work, we addressed the limitations of previous photography composition methods by introducing photography perspective composition (PPC), a novel paradigm that extends beyond 2D cropping to achieve 3D recomposition. Our approach is inspired by real-world street photography practices where photographers use perspective adjustment to establish better relative relationships between subjects. To overcome the challenges of implementing PPC, particularly the lack of suitable datasets and unclear assessment criteria, we made three significant contributions. We developed a framework for automatically constructing a PPC dataset from expert photographs, created a system for generating perspective transformation videos that guide users from less favorable to aesthetically enhanced views, and introduced the perspective quality assessment (PQA) model that evaluates both video quality and compositional aesthetics. We hope this work opens up new possibilities in computational photography and inspires further research in perspective-aware composition.
@@
Figure 11: Diverse data presents instability.
(3) Data Scaling Behavior. Despite the strong scene diversity provided by numerous expert photography images, we observed that model outputs become unstable as the training data volume increases. As shown in Tab. 3c, Fig. 4a, and Fig. 11, while both accuracy and quality initially improve with increasing data volume, performance deteriorates when the training set size grows further. We hypothesize that this challenge lies in maintaining the desired model behavior to ensure proper perspective rather than deviating into unintended random behaviors. This phenomenon bears similarities to what was described in IC-Light zhang2025scaling , and we plan to incorporate their training methodology to investigate whether it can enhance training stability in our future work.